Electrons are emitted with zero velocity from a certain metal surface when it is exposed to radiations of lamda=6800A. Calculate threshold frequency and work function of metal.


1
Expert's answer
2021-02-16T15:56:39-0500

Let's use the Einstein equation for the photoelectric effect

"h\nu=A+E_k"

Where

"A-" Electron work function

"E_k-" Kinetic energy of an electron

By the condition of the problem

"E_k=0"

Then we write

"A=h\nu=\frac{hc}{\lambda}=\frac{6.63 \cdot 10^{-34}J \cdot s \cdot 3 \cdot 10^{8} \cdot m/s}{6.8 \cdot 10^{-7} m}=2.925 \cdot 10^{-19}J"

"\nu_{min}=\frac{A}{h}=\frac{2.925 \cdot 10^{-19}}{6.63 \cdot 10^{-34}}=4.41 \cdot 10^{14} \cdot \frac {1}{s}"